### v1

# Construction of Solutions with Extraordinary Gradient Amplification and Localization for Schrödinger Equations

Abstract.
This paper constructs solutions to linear and nonlinear Schrödinger-type equations in two and three spatial dimensions that exhibit prescribed, extraordinary gradient amplification and localization. For any finite time interval $[0,T]$, any prescribed collection of $n\in\mathbb{N}$ distinct points on $\partial D$, where $D$ is the compact support of the anisotropic coefficients, lower-order terms, or nonlinearities, and any amplitude threshold $\mathcal{M}>0$, we show that one can design smooth initial and/or boundary data such that the spatial gradients of the resulting solutions exceed $\mathcal{M}$ in neighborhoods of these points outside $D$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\frac{1}{2}}$-norm of the solution near each prescribed point outside $D$ and the $C^{1,\frac{1}{2}}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$ for almost every $t\in[0,T]$.
We further prove that the spatial measure of the regions where the gradient magnitude exceeds $\mathcal{M}$ tends to zero as $\mathcal{M}\to\infty$, demonstrating that the amplification phenomenon is highly localized. This effect arises from the structure of the Schrödinger-type equation combined with carefully designed input profiles. From a physical perspective, the results provide a deterministic analogue of localization phenomena observed in quantum scattering and Anderson localization. In addition, the observed trade-off between extreme spatial localization and large gradient amplification is fully consistent with the spirit of the Heisenberg uncertainty principle: while the latter is traditionally formulated in a global $L^{2}$ space–frequency framework, our results offer a complementary deterministic manifestation at the level of localized spatial gradients in Schrödinger dynamics.
Keywords:  Schrödinger equation; localized gradient amplification; localized surface resonance; customized Anderson localization; quantum wave scattering; Heisenberg uncertainty principle
2020 Mathematics Subject Classification:  35Q55,35B65,35P25,35B40,47A75,35B30,81Q10
Huaian Diao
School of Mathematics and Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education, Jilin University, Changchun 130012, China
E-mail: o@jlu.edu.cn, hadiao@gmail.com
Xieling Fan
School of Mathematics, Central South University, Changsha, China and Department of Mathematics, City University of Hong Kong, Kowloon, Hong Kong SAR, China
E-mail: fanxieling@outlook.com, xielinfan2-c@my.cityu.edu.hk
Hongyu Liu
Department of Mathematics, City University of Hong Kong, Kowloon, Hong Kong SAR, China
E-mail: hongyu.liuip@gmail.com, hongyliu@cityu.edu.hk
1. Introduction
1.1. Mathematical setup and summary of major results
Initially we focus on the mathematical models for both linear and nonlinear Schrödinger equations in two and three spatial dimensions in this paper, where the mechanism for these mathematical models will be elaborated in more detail later. Let the unknown function $u:\Omega\times[0,T]\rightarrow\mathbb{C}$ satisfy the homogeneous initial-boundary value problem for a linear Schrödinger equation:
$\begin{cases}\mathrm{i}\partial_{t}u+L_{1}u=0&\text{in }\Omega\times(0,T],\\ u(\mathbf{x},0)=\phi(\mathbf{x})&\text{in }\Omega,\\ u(\mathbf{x},t)=\psi(\mathbf{x},t)&\text{on }\partial\Omega\times[0,T],\end{cases}$ (1.1)
and let $U:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{C}$ satisfy the Cauchy problem for a nonlinear Schrödinger equation:
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}U+L_{2}U+N(U)=0&&\text{in }\mathbb{R}^{d}\times(0,T],\\ &U(\mathbf{x},0)=\Phi(\mathbf{x})&&\text{in }\mathbb{R}^{d},\end{aligned}\right$. (1.2)
where the following conditions hold:
  1. (1)
    $\mathrm{i}$ denotes $\sqrt{-1}$, $\Omega$ is a bounded Lipschitz domain in $\mathbb{R}^{d}$, $d=2,3$, and $0<T<\infty$.
  2. (2)
    $L_{1}=\nabla\cdot(\mathbf{A}_{1}\nabla)$, $L_{2}=\nabla\cdot(\mathbf{A}_{2}\nabla)+c$.
  3. (3)
    The complex-valued coefficient $\mathbf{A}_{1}=(a_{1}^{ij}(\mathbf{x}))_{i,j=1}^{d}$ satisfies:
    1. (a)
      $\overline{\mathbf{A}}_{1}^{\top}\!(\mathbf{x})=\mathbf{A}_{1}(\mathbf{x}),~\forall\mathbf{x}\in\Omega$;
    2. (b)
      $\overline{\xi}^{\top}\mathbf{A}_{1}(\mathbf{x})\xi\geq\theta\lvert\xi\rvert^{2}$ for a.e. $\mathbf{x}\in\Omega$, $\forall\xi\in\mathbb{C}^{d}$, with constant $\theta>0$;
    3. (c)
      $\mathbf{A}_{1}\in W^{2,\infty}(\Omega)\cap C^{1,1}(\overline{\Omega})$ ;
    where $\overline{\cdot}$ denotes the complex conjugate and $\top$ denotes the transpose.
  4. (4)
    $D$ is a bounded Lipschitz domain such that $D\Subset\Omega$, $\D¯\Omega\backslash\overline${D} is connected and $\mathbf{A}_{1}$ has compact support in $D$, i.e.,
    $\operatorname{supp}(\mathbf{A}_{1}-\mathbf{I})\subset D.$
  5. (5)
    The complex-valued coefficients $\mathbf{A}_{2}(\mathbf{x},t)=\big(a_{2}^{ij}(\mathbf{x},t)\big)_{i,j=1}^{d},c(\mathbf{x},t)$ satisfy the following conditions:
    1. (a)
      The matrix $\mathbf{A}_{2}(\mathbf{x},t)$ is given by
      $\mathbf{A}_{2}(\mathbf{x},t)=\begin{cases}\mathbf{B}(t),&(\mathbf{x},t)\in\overline{D}\times[0,T],\\[4.0pt] \mathbf{I},&(\mathbf{x},t)\in(\mathbb{R}^{d}\setminus\overline{D})\times[0,T],\end{cases}$
      where $\mathbf{I}$ is the $d$-dimensional identity matrix and $\mathbf{B}(t)=\big(b^{ij}(t)\big)_{i,j=1}^{d}$ is a Hermitian matrix and $\mathbf{B}\in L^{\infty}[0,T]$. Moreover, it satisfies the uniform ellipticity condition
      $b^{ij}(t)\,\xi_{i}\,\overline{\xi_{j}}\geq\theta|\xi|^{2},\quad\forall\,\xi\in\mathbb{C}^{d},\;\forall\,t\in[0,T].$
    2. (b)
      The function $c\in L^{\infty}\bigl(0,T;W^{3,\infty}(\mathbb{R}^{d})\bigr),$ and its spatial support is contained in $D$ for all $t\in[0,T]$, i.e.,
      $\operatorname{supp}c(\cdot,t)\subset D,\qquad\forall t\in[0,T].$
  6. (6)
    The nonlinearity $N:\mathbb{C}\to\mathbb{C}$ is a complex polynomial given by
    $N(U)=\sum_{k=2}^{l_{0}}\alpha_{k}U^{k},$ (1.3)
    where $l_{0}\geq 2$ is a fixed integer. The coefficients $\alpha_{k}$ satisfy::
    $\alpha_{k}\in C^{\infty}([0,\infty)\times\mathbb{R}^{d};\mathbb{C}),\quad k=2,\dots,l_{0},$
    with compact support in $D$, i.e.,
    $\operatorname{supp}(\alpha_{k}(\cdot,t))\subset D,\quad\forall t\in[0,\infty).$
  7. (7)
    The initial data $\phi\in C^{\infty}{(\Omega)}$, $\Phi\in C^{\infty}(\mathbb{R}^{d})$, and boundary data $\psi\in C^{\infty}([0,T];C(\partial\Omega))$ will be specified in Section 3.
In this paper, we identify and construct a mechanism that produces highly localized gradient amplification of solutions to both the linear and nonlinear Schrödinger-type equations in two and three spatial dimensions. The phenomenon is described by the following two features:
  1. (i)
    For any prescribed set of $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, where $D$ is the support of the anisotropic coefficients $\mathbf{A}_{1}$ in (1.1) or $\mathbf{A}_{2}$ in (1.2), the zero-order term $c$ in (1.2), and the nonlinear term $N$ in (1.2), we rigorously prove—by selecting suitably tuned smooth initial–boundary data for (1.1) or smooth initial data for (1.2)—that there exist localized gradient amplification regions in neighborhoods of each $\mathbf{x}_{i}$ located outside $D$ ($i=1,\ldots,n$). In these regions, the $C^{0}$-norms of the spatial gradients of the solutions, namely $\nabla u(t)$ to (1.1) and $\nabla U(t)$ to (1.2), exceed any prescribed large constant $\mathcal{M}$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\alpha}$-norm of the solution in these regions and the $C^{1,\alpha}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$. These properties are established in Theorems 1.1 and 1.2, where the time horizon $T\in\mathbb{R}_{+}$ is arbitrary.
  2. (ii)
    We further provide precise quantitative estimates for the measure of the localized gradient amplification regions in terms of the amplitude threshold $\mathcal{M}$, which characterizes the strength of the gradient amplification for the solutions $u$ and $U$ to (1.1) and (1.2), respectively. Our analysis shows that, as $\mathcal{M}\to\infty$, the measure of these regions tends to zero, demonstrating that the gradient amplification phenomenon is sharply localized in space.
We are now in a position to state our main theorems. For initial-boundary value problem of a linear Schrödinger-type equation (1.1), we have:
Theorem 1.1.
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial and boundary inputs ((as specified in (3.21))), there exists a small parameter
$r_{0}=\min\left\{\,\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|,\ \frac{1}{3}\operatorname{dist}(D,\partial\Omega)\,\right\}$
such that problem (1.1) admits a unique solution $u$ satisfying ::
$u\in L^{\infty}\big([0,T];H^{3}_{\mathrm{loc}}(\Omega)\big),$
with the following properties ::
  1. (1)
    Interior improved Hölder regularity :: for any bounded Lipschitz domain $\Omega^{\prime}\Subset\Omega$,
    $u\in L^{\infty}\big([0,T];C^{1,\frac{1}{2}}(\overline{\Omega}^{\prime})\big).$
  2. (2)
    Simultaneous localized gradient amplification near the points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}$ in the following sense: for almost every $t\in[0,T]$ and for each $i=1,2,\ldots,n$,
    $\|\nabla u(\cdot,t)\|_{C^{0}\!\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right)}\geq\mathcal{M},$
    and
    $\frac{\|u(\cdot,t)\|_{C^{1,\frac{1}{2}}\!\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right)}}{\|u(\cdot,t)\|_{C^{1,\frac{1}{2}}(\overline{D})}}>\frac{\mathcal{M}}{2}.$
  3. (3)
    Vanishing measure of the localized gradient amplification set :: for almost every $t\in[0,T]$,
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla u(\mathbf{x},t)|>\mathcal{M}\right\}$
[…]
    as $\mathcal{M}\to\infty$.
Similarly, for Cauchy problem of a nonlinear Schrödinger-type equation (1.2), we have:
Theorem 1.2.
For any arbitrarily large $\mathcal{M}>0$, $n$ distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, and any time interval $[0,T]$ with $0<T<\infty$, by prescribing appropriate initial inputs ((as specified in (3.22))), there exists a small parameter
$r_{0}=\min\left\{\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|\right\}$
such that problem (1.2) admits a unique solution $U$ satisfying:
$U\in L^{\infty}\big(0,T;H^{3}_{\mathrm{loc}}(\mathbb{R}^{d})\big),$
with the following properties::
  1. (1)
    Hölder regularity: For any bounded Lipschitz domain $\Omega^{\prime}\subset\mathbb{R}^{d}$, U L^([0,T]; C^1,12(Ω’)).
  2. (2)
    Simultaneous localized gradient amplification near the points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}$ in the following sense: for almost every $t\in[0,T]$ and for each $i=1,2,\ldots,n$,
    $\|\nabla U(\cdot,t)\|_{C^{0}\!\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right)}\geq\mathcal{M},$
    and
    $\frac{\|U(\cdot,t)\|_{C^{1,\frac{1}{2}}\!\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right)}}{\|U(\cdot,t)\|_{C^{1,\frac{1}{2}}(\overline{D})}}>\frac{\mathcal{M}}{2}.$
  3. (3)
    Vanishing measure of the localized gradient amplification set: for almost every $t\in[0,T]$,
    $\operatorname{meas}\left\{\mathbf{x}\in\bigcup_{i=1}^{n}\left(B_{3r_{0}}(\mathbf{x}_{i})\setminus\overline{D}\right):|\nabla U(\mathbf{x},t)|>\mathcal{M}\right\}$
[…]
    as $\mathcal{M}\to\infty$.
The construction mechanism developed in this work for producing customized solutions to Schrödinger equations with localized gradient amplification exhibits the following key features.
First, it applies to a broad class of linear and nonlinear Schrödinger-type equations in two and three spatial dimensions. Inside the region $D$, the coefficients may be quite general (e.g., smooth anisotropic coefficients or nonlinearities given by arbitrary polynomials), while outside $D$ the equation coincides with the standard Schrödinger equation.
Second, for any prescribed time horizon $T\in\mathbb{R}_{+}$, any arbitrarily large amplification threshold $\mathcal{M}>0$, and any finite collection of distinct points $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}\in\partial D$, one can construct suitably designed smooth initial and/or boundary data such that, for both the linear problem (1.1) and the nonlinear problem (1.2), the magnitude of the spatial gradient of the corresponding solution exceeds $\mathcal{M}$ in neighborhoods of each prescribed point located outside $D$ for almost every $t\in[0,T]$. Moreover, the ratio between the local $C^{1,\alpha}$-norm of the solution in these neighborhoods and the $C^{1,\alpha}$-norm inside $D$ is bounded from below by $\mathcal{M}/2$. The time interval $T$ can be chosen arbitrarily.
Third, the regions where the gradient amplification occurs are contained in the balls $B_{3r_{0}}(\mathbf{x}_{i})$. As the aforementioned bound $\mathcal{M}\to\infty$, the Lebesgue measure of these regions tends to zero, in agreement with the well-posedness of the underlying problems. This illustrates a clear trade-off: achieving larger gradient magnitude requires stronger spatial concentration, whereas milder amplification allows the region to occupy a larger set. Moreover, when the points $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ lie close to one another, the corresponding amplification regions must remain sufficiently confined to avoid interference.
To the best of our knowledge, this work provides the first systematic and customized construction of solutions exhibiting localized gradient amplification for Schrödinger-type equations in both linear and nonlinear regimes. A more detailed discussion of the underlying ideas and implications is given in Subsection 1.3. We conclude this subsection with several remarks concerning the methodology employed in the proofs and the relationship between the assumptions and conclusions of Theorems 1.1 and 1.2.
Remark 1.3.
The localized gradient amplification mechanism developed in this work for Schrödinger equations is rooted in spectral analysis, in particular the use of transmission eigenfunctions associated with the interior transmission problem [5] and their approximation by Herglotz wave functions. By exploiting these tools, we construct a carefully designed auxiliary function $u_{0}$ (see Theorem 3.1), which serves as the fundamental building block for prescribing the initial and boundary inputs in both the linear and nonlinear Schrödinger equations. Further details of this construction are provided in Section 3.
Remark 1.4.
Both the solution $u(t)$ to the linear problem (1.1) and the solution $U(t)$ to the nonlinear problem (1.2) exhibit local $H^{3}$-regularity for almost every $t\in[0,T]$. In the linear case, this regularity follows from the interior elliptic regularity theorem applied to $u(t)$. In the nonlinear setting, the regularity of $U(t)$ is ensured by the smoothness assumptions on the coefficients together with a Banach fixed point argument. Consequently, in both cases, the Sobolev embedding theorem implies that the solutions possess $C^{1,\frac{1}{2}}$-regularity in neighborhoods of the localized gradient amplification points for almost every $t\in[0,T]$, which guarantees that the spatial gradients are well defined in the classical sense. Moreover, if additional time-regularity assumptions are imposed on the coefficients, the above regularity results can be strengthened to hold for all $t\in[0,T]$.
Remark 1.5.
In this work, the localized gradient amplification regions are described using balls in 3D and disks in 2D. It is plausible that other types of sets could be employed to characterize such highly localized gradient regions, although doing so would require more refined technical arguments.
Remark 1.6.
For both the linear problem (1.1) and the nonlinear problem (1.2), the time $T\in\mathbb{R}_{+}$ of the interval $[0,T]$ can be chosen arbitrarily large while still guaranteeing the localized gradient amplification of the solutions. In the linear case, this follows directly from the classical well-posedness theory for linear Schrödinger equations. In the nonlinear case, however, such long-time existence requires a more refined analysis. By carefully tuning the initial data and combining spectral analysis with an application of the Banach fixed-point theorem, we are able to extend the solution to (1.2) over any prescribed time interval $T$. For further details, see Remark 5.3 and the proof of Theorem 1.2.
1.2. Existing studies and motivation
We begin by recalling several classical results on singularity formation in the nonlinear Schrödinger equation
$iu_{t}=\Delta u+\mu|u|^{p-1}u,\qquad x\in\mathbb{R}^{d},\ t>0,\qquad u(x,0)=u_{0}(x).$
Blow-up refers to the existence of a finite time $T>0$ such that a suitable norm of the solution $u(t)$ becomes unbounded as $t\to T$. In the $L^{2}$-supercritical case $p>1+4/d$, Glassey [11] showed that for the focusing equation $(\mu=-1)$, smooth initial data with sufficient decay at infinity and satisfying certain energy and momentum conditions lead to finite-time blow-up. In the $L^{2}$-subcritical case $1<p<1+4/d$, global well-posedness holds for the defocusing equation $(\mu=1)$ for all finite-energy data, whereas in the focusing case $(\mu=-1)$ global existence is guaranteed below specific mass-energy thresholds [10]. In the $L^{2}$-critical case $p=1+4/d$, Merle [16] demonstrated that solutions can blow up at any prescribed finite set of points, with the blow-up profile concentrating around these locations while vanishing elsewhere. Further refinements of multi-point blow-up, including log–log type concentration, are given in [8]. These results highlight the remarkable flexibility of singularity formation in the critical regime, where spatial localization can be prescribed precisely. For general background on the nonlinear Schrödinger equation, see [20, 19, 4].
In the present work, we establish an analogous, yet fundamentally different, phenomenon for certain linear and nonlinear Schrödinger equations, where solutions remain local $H^{3}$-regular over the entire interval $[0,T]$ but nevertheless develop sharply localized regions of high spatial gradient at arbitrarily selected points. This controlled formation of high-gradient regions draws conceptual inspiration from localization phenomena extensively studied in both mathematical and physical Schrödinger theory. In systems with random or quasi-periodic potentials, Anderson localization [2, 6] demonstrates that eigenfunctions may concentrate strongly around specific spatial locations even in the absence of nonlinear effects. Such behavior plays a fundamental role in quantum scattering, wave propagation in disordered media, and the spectral analysis of Schrödinger operators. More generally, classical quantum scattering theory shows that potential landscapes, boundary conditions, and geometric trapping can generate highly localized wave functions, resonances, or tunneling effects [21, 24, 17], which govern transport and energy distribution in quantum systems.
From a physical perspective, the localization mechanism uncovered in this work is also consistent with the qualitative content of the Heisenberg uncertainty principle [12]. The uncertainty principle asserts a fundamental trade-off between spatial localization and frequency (or momentum) concentration, typically formulated in a global $L^{2}$-based phase-space framework. In our setting, although no probabilistic or Fourier-analytic uncertainty inequality is invoked, an analogous trade-off emerges at the level of deterministic PDE dynamics: as the spatial gradients of the solution become arbitrarily large near prescribed points, the regions in which this amplification occurs necessarily shrink in measure. Thus, extreme spatial localization is accompanied by increasingly large gradients, reflecting a complementary manifestation of uncertainty-type behavior within Schrödinger-type evolutions. Our results therefore provide a rigorous and deterministic illustration of how localization and high-frequency effects can coexist in regular solutions, bridging ideas from PDE analysis, quantum scattering, and foundational principles of quantum mechanics.
1.3. Main contributions and technical discussions
Our work constructs novel solutions for a broad class of Schrödinger-type equations defined by their behavior inside and outside a bounded domain $D$. Specifically, we consider Schrödinger equations that reduce to the standard Schrödinger equation outside $D$, while inside $D$ they are governed by variable-coefficient operators and may include arbitrary polynomial nonlinearities. These solutions with localized gradient amplification behavior exhibit the following precise characteristics:
  1. (i)
    For any prescribed $\mathcal{M}>0$ and any prescribed number $n$ of distinct points $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ on $\partial D$, we construct inputs (specifically, initial and/or boundary inputs) that yield solutions $u$ and $U$ to problems (1.1) and (1.2), respectively. In the specified regions near these points, the $L^{\infty}$-norms of $\nabla u$ and $\nabla U$ exhibit localized gradient amplification bounded below by $\mathcal{M}$ for almost every time $t$ in the interval $[0,T]$, where $T$ can be chosen arbitrarily in advance. Moreover, the amplification is quantitatively contrasted with the behavior inside $D$: the ratio between the local $C^{1,1/2}$-norm of the solution in each amplification region outside $D$ and the $C^{1,1/2}$-norm of the solution restricted to $\overline{D}$ is bounded from below by $\mathcal{M}/2$ for almost every $t\in[0,T]$. This demonstrates that the observed amplification is not merely absolute, but also relative to the interior regularity of the solution. For the nonlinear problem (1.2), the solution exists for an arbitrarily large time interval. This long-time existence is achieved through a combination of spectral analysis, application of the Banach fixed-point theorem, and careful tuning of the initial data (see Remark 5.3 and the proof of Theorem 1.2).
  2. (ii)
    Due to the well-posedness of problems (1.1) and (1.2), the measure of the localized gradient amplification regions—where the gradient of solutions is bounded below by $\mathcal{M}$—vanishes for almost every $t\in[0,T]$ as $\mathcal{M}\to\infty$. We provide detailed characterizations of the measure of these regions in Theorems 1.1 and 1.2. Stronger localized gradient amplification corresponds to smaller spatial support of these regions. This quantitative decay property of the measure with respect to $\mathcal{M}$ contrasts with classical blowup, where the singular set typically has positive measure at blowup time. Importantly, under the specified smooth and finely tuned inputs—even with the weaker regularity coefficients specified in Items 3 and 5 of (1.1) or (1.2) whose supports lie in a bounded Lipschitz domain—the solutions $u$ and $U$ retain local spatial $H^{3}$ regularity for almost every $t\in[0,T]$. Moreover, within the localized gradient amplification regions, the solutions maintain local spatial $C^{1,1/2}$ regularity for almost every $t\in[0,T]$.
This localized gradient amplification behavior is realized through a careful combination of spectral theory for transmission eigenfunctions of the interior transmission problem [5] and Herglotz wave function approximations. By analyzing the spectral properties of these eigenfunctions, we identify key structures that allow for the selective amplification of gradients near pre-specified points. Herglotz approximations are then employed to explicitly construct smooth initial and/or boundary inputs that excite these structures, effectively transferring the amplification to the solutions of both the linear problem (1.1) and the nonlinear problem (1.2). The resulting inputs, detailed in (3.21) and (3.22), guarantee that the gradient of the solution attains the prescribed bound $\mathcal{M}$ in the targeted regions, while maintaining regularity.
The remainder of this paper is organized as follows. Section 2 establishes the requisite mathematical preliminaries. In Section 3, we develop a unified framework for constructing initial and boundary inputs for both problems (1.1) and (1.2). Section 4 presents the complete proof of Theorem 3.1, establishing well-posedness for the linear auxiliary problem (4.1) and constructing solutions with localized gradient amplification through strategic parameter tuning. Section 5 extends this methodology to the nonlinear setting in Theorem 1.2, where a fixed-point argument establishes well-posedness for the auxiliary problem (5.1) and proves the analogous localized gradient amplification behavior of the solution.
2. Preliminaries
Throughout this paper, the constant $C$ may change from line to line. In the statements of theorems, we explicitly specify the parameters on which constants depend. However, within proofs, for brevity and to avoid cumbersome notation, we typically do not detail the precise dependence of constants on various parameters. For particularly significant constants that play a crucial role in the argument, we assign specific labels such as $C_{1}$, $C_{2}$, and so forth. When the dependence on certain key parameters is essential to the analysis, we may highlight this relationship through notation such as $C_{2}(r_{0})$ to indicate that the constant $C_{2}$ depends specifically on the parameter $r_{0}$.
We adopt the following notation:
$\nabla u=\nabla_{\mathbf{x}}u=\left(\partial_{x_{1}}u,\partial_{x_{2}}u,\dots,\partial_{x_{d}}u\right)=\left(u_{x_{1}},u_{x_{2}},\dots,u_{x_{d}}\right),\partial_{t}u=u_{t}=u^{\prime},\partial_{t}^{2}u=u_{tt}=u^{\prime\prime},$
and similarly for higher-order derivatives. The symbols $\Re u$ and $\Im u$ denote the real and imaginary parts of $u$, respectively, and $\mathrm{i}$ denotes the imaginary unit $\sqrt{-1}$. The floor function is written as $\lfloor x\rfloor$.
For an open set $U\subseteq\mathbb{R}^{n}$, an integer $k\geq 0$, and $0<\gamma\leq 1$, the Hölder space $C^{k,\gamma}(\bar{U})$ consists of all functions $u\in C^{k}(\bar{U})$ for which the norm
$\|u\|_{C^{k,\gamma}(\bar{U})}=\sum_{|\alpha|\leq k}\sup_{\bar{U}}|D^{\alpha}u|+\sum_{|\alpha|=k}[D^{\alpha}u]_{C^{\gamma}(\bar{U})}$
is finite, where $v]_{C^{\gamma}(\bar{U})}$ denotes the $\gamma$-Hölder seminorm:
$v]_{C^{\gamma}(\bar{U})}=\sup_{\begin{subarray}{c}\mathbf{x},\mathbf{y}\in U\\ \mathbf{x}\neq\mathbf{y}\end{subarray}}\frac{|v(\mathbf{x})-v(\mathbf{y})|}{|\mathbf{x}-\mathbf{y}|^{\gamma}}.$
Let $1\leq p\leq\infty$ and let $X$ be a Banach space. The Bochner–Sobolev space $W^{1,p}(0,T;X)$ consists of all functions $u\in L^{p}(0,T;X)$ such that the first weak derivative $u^{\prime}$ exists and belongs to $L^{p}(0,T;X)$. The norm on $W^{1,p}(0,T;X)$ is defined as
$\|u\|_{W^{1,p}(0,T;X)}=\begin{cases}\left(\int_{0}^{T}\|u(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime}(t)\|_{X}^{p}dt\right)^{\frac{1}{p}}&\text{if }1\leq p<\infty,\\ \operatorname{ess\,sup}_{0\leq t\leq T}\left(\|u(t)\|_{X}+\|u^{\prime}(t)\|_{X}\right)&\text{if }p=\infty.\end{cases}$
Similarly, the Bochner Sobolev space $W^{2,p}(0,T;X)$ consists of all functions $u\in L^{p}(0,T;X)$ such that both the first and second weak derivatives $u^{\prime}$ and $u^{\prime\prime}$ exist and belong to $L^{p}(0,T;X)$. The norm on $W^{2,p}(0,T;X)$ is given by
$\|u\|_{W^{2,p}(0,T;X)}=\begin{cases}\left(\int_{0}^{T}\|u(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime}(t)\|_{X}^{p}dt+\int_{0}^{T}\|u^{\prime\prime}(t)\|_{X}^{p}dt\right)^{\frac{1}{p}}&\text{if }1\leq p<\infty,\\ \operatorname{ess\,sup}_{0\leq t\leq T}\left(\|u(t)\|_{X}+\|u^{\prime}(t)\|_{X}+\|u^{\prime\prime}(t)\|_{X}\right)&\text{if }p=\infty.\end{cases}$
Equipped with these norms, both $W^{1,p}(0,T;X)$ and $W^{2,p}(0,T;X)$ are Banach spaces.
Let $J_{m}$ denote the Bessel function of the first kind and $j_{m}$ the spherical Bessel function. Let $j_{m,s}$ denote the $s$-th positive zero of $J_{m}$ and $j_{m,s}^{\prime}$ the $s$-th positive zero of $J_{m}^{\prime}$. According to [1] and [14], the following properties hold:
$m\leq j_{m,1}^{\prime}<j_{m,1}<j_{m,2}^{\prime}<j_{m,2}<j_{m,3}^{\prime}<\cdots.$
The Bessel function admits the Weierstrass factorization:
$J_{m}(x)=\frac{(|x|/2)^{m}}{\Gamma(m+1)}\prod_{s=1}^{\infty}\left(1-\frac{x^{2}}{j_{m,s}^{2}}\right).$
For integer or positive $\alpha$, the Bessel function has the power series expansion:
$J_{\alpha}(x)=\sum_{m=0}^{\infty}\frac{(-1)^{m}}{m!\,\Gamma(m+\alpha+1)}\left(\frac{x}{2}\right)^{2m+\alpha},\quad x>0,$ (2.1)
where $\Gamma(x)$ is the gamma function. The relationship between $J_{m}$ and $j_{m}$ is given by:
$j_{m}(x)=\sqrt{\frac{\pi}{2x}}J_{m+\frac{1}{2}}(x).$ (2.2)
The spherical harmonics $Y_{m}^{l}(\theta,\varphi)$ (see [5, Theorem 2.8]) are defined as:
$Y_{m}^{l}(\theta,\varphi)=\sqrt{\frac{2m+1}{4\pi}\frac{(m-|l|)!}{(m+|l|)!}}P_{m}^{|l|}(\cos\theta)e^{im\varphi},$ (2.3)
where $m=0,1,2,\ldots$, $l=-m,\ldots,m$, and $P_{m}^{|l|}$ are the associated Legendre functions. The following estimate for Legendre functions will be used in the subsequent analysis.
Lemma 2.1.
(see [15], Corollary 3). For real $x\in[-1,1]$ and integers $m$, $n$ with $1\leq m\leq n$,
$\frac{1}{\sqrt{2.22(m+1)}}<\max_{x\in[-1,1]}|P_{n}^{m}(x)|\sqrt{\frac{(n-m)!}{(n+m)!}}<\frac{2^{5/4}}{\pi^{3/4}}\frac{1}{m^{1/4}}.$
In what follows, we assume that $\omega$ is a fixed constant. The Herglotz function is defined as
$H_{g}(\mathbf{x})=\int_{\mathbb{S}^{d-1}}g(\theta)\exp(\mathrm{i}\omega\mathbf{x}\cdot\theta)d\theta,\quad\mathbf{x}\in\mathbb{R}^{d},$ (2.4)
where $g\in L^{2}(\mathbb{S}^{d-1})$ is the Herglotz kernel. We now state a key lemma that will be used in the next section.
Lemma 2.2.
(see [23]) Let $\mathcal{O}$ be a bounded domain of class $C^{\alpha,1}$, $\alpha\in\mathbb{N}\cup\{0\}$, in $\mathbb{R}^{d}$. Denote by $\mathbb{H}$ the space of all Herglotz functions of the form (2.4). Define
$\mathbb{H}(\mathcal{O}):=\left\{\left.u\right|_{\mathcal{O}}:u\in\mathbb{H}\right\}$
and
$\mathfrak{H}(\mathcal{O}):=\left\{u\in C^{\infty}(\mathcal{O}):\Delta u+\omega^{2}u=0\text{ in }\mathcal{O}\right\}.$
Then $\mathbb{H}(\mathcal{O})$ is dense in $\mathfrak{H}(\mathcal{O})\cap H^{\alpha+1}(\mathcal{O})$ with respect to the $H^{\alpha+1}(\mathcal{O})$-norm.
Remark 2.3.
The Herglotz function $H_{g}$ defined in (2.4) satisfies the Helmholtz equation $\Delta H_{g}+\omega^{2}H_{g}=0$ in $\mathbb{R}^{d}$. By standard elliptic regularity theory, $H_{g}$ is analytic with respect to the spatial variable. A key property of Herglotz functions is that they are generally not square-integrable over the entire space; specifically, $H_{g}\in L^{2}_{\text{loc}}(\mathbb{R}^{d})$ but $H_{g}\notin L^{2}(\mathbb{R}^{d})$ for non-trivial kernels $g$.
3. A general framework for specifying inputs
In this section, we establish the general framework for constructing initial and boundary inputs for problems (1.1) and (1.2). To achieve this, we employ spectral analysis, interior transmission problems, and Herglotz approximation techniques to construct a function $u_{0}$ that will serve as the foundation for our desired inputs.
[…]
Figure 1. Geometric configuration where each external point $\mathbf{y}_{i}$ is placed at a uniform distance from its corresponding boundary point $\mathbf{x}_{i}\in\partial D$.
Theorem 3.1.
For any predetermined positive parameters $\varepsilon<1$, $r_{0}$ sufficiently small, and $n$ distinct points $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\partial D$, we can properly choose corresponding $n$ balls satisfying:
  1. (i)
    For problem (1.1) ((bounded domain $\Omega$ with $D\Subset\Omega):$ $B_{r_{0}}(\mathbf{y}_{i})\Subset\Omega\setminus\overline{D}$,
  2. (ii)
    For problem (1.2) ((full space $\mathbb{R}^{d}):$ $B_{r_{0}}(\mathbf{y}_{i})\Subset\mathbb{R}^{d}\setminus\overline{D}$,
such that $|\mathbf{x}_{i}-\mathbf{y}_{i}|=2r_{0}$, the balls are mutually disjoint and
$\operatorname{dist}(\partial D,B_{r_{0}}(\mathbf{y}_{i}))>\frac{r_{0}}{4},\ \forall i=1,2...,n.$
Then, there exists an integer $M:=M(r_{0})$ and a smooth function $u_{0}$ with the following properties:
  1. (1)
    $u_{0}$ has the form::
    $u_{0}(\mathbf{x},t)=H_{g}(\mathbf{x})\exp(-\mathrm{i}t),$
    where $H_{g}(\mathbf{x})$ is a Herglotz function given by (3.15).
  2. (2)
    $u_{0}$ satisfies the free Schrödinger equation::
    $\mathrm{i}\partial_{t}u_{0}+\Delta u_{0}=0\quad\text{in }\mathbb{R}^{d}\times[0,\infty).$
  3. (3)
    It exhibits small $C^{1,\frac{1}{2}}$-norm in $\overline{D}:$
    $u_{0}(\cdot,t)||_{C^{1,\frac{1}{2}}(\overline{D})}<C_{1}\varepsilon\quad\text{for all}~t\in[0,\infty),$ (3.1)
    where constant $C_{1}$ depends on $D$ and $d$.
  4. (4)
    For any predetermined integer $m\geqslant M(r_{0},\omega)$, there exist points $\mathbf{y}^{*}_{i}\in\partial B_{r_{0}}(\mathbf{y}_{i})$ ($i=1,\dots,n$) with large amplitude $u_{0}(\mathbf{y}^{*}_{i},t)$ such that
    $|u_{0}(\mathbf{y}^{*}_{i},t)|$ $>\frac{\sqrt{2m+2}\,r_{0}^{-1}}{2\sqrt{2\pi}}-C_{2}\varepsilon,$ $d$ $=2,$ (3.2)
    $|u_{0}(\mathbf{y}^{*}_{i},t)|$ $>\frac{\sqrt{2m+3}\,r_{0}^{-3/2}}{16}-C_{2}\varepsilon,$ $d$ $=3,$ (3.3)
    for all $t\in[0,\infty)$, where constant $C_{2}$ depends on $r_{0}$, $D$, and $d$.
Remark 3.2.
From now and remaining of this paper, we denote by $\varepsilon$, $r_{0}$ the parameters, and by $u_{0}$ the function, specified in Theorem 3.1.
Remark 3.3.
For notation convenience, we denote $C_{2}:=C_{2}(r_{0})$. In particular, by (3.17), (3.20), and the Sobolev embedding theorem, $C_{2}\to\infty$ as $r_{0}\to 0$. The constants $C_{1}$ and $C_{2}$ will be used to determine $\varepsilon$, $r_{0}$, and $m$ in the proofs of Theorems 1.1 and 1.2.
Remark 3.4.
The construction distinguishes between two physical settings:
  1. (i)
    For the linear problem (1.1) in the bounded domain $\Omega$ (with $D\Subset\Omega$), the balls satisfy $B_{r_{0}}(\mathbf{y}_{i})\Subset\Omega\setminus\overline{D}$.
  2. (ii)
    For the nonlinear problem (1.2) in the full space $\mathbb{R}^{d}$, the balls satisfy $B_{r_{0}}(\mathbf{y}_{i})\Subset\mathbb{R}^{d}\setminus\overline{D}$.
Importantly, the selection of these balls is highly flexible they can be chosen arbitrarily within the respective domains, subject only to the above containment conditions. Despite these domain differences, the constructed $u_{0}$ shares identical properties in both cases. For simplicity, we do not distinguish between the corresponding $u_{0}$ functions in subsequent analysis.
Prior to proving Theorem 3.1, we establish the following auxiliary result:
Lemma 3.5.
For any fixed $r_{0}$, the series $I_{i}(m,r_{0})\rightarrow 0$ as $m\rightarrow+\infty$ for $i=1,2,3,4$, where
$I_{1}(m,r_{0})$ $:=\sum_{k=1}^{\infty}\frac{(-1)^{k}\Gamma(m+\frac{3}{2})\sqrt{2m+3}}{k!\Gamma(m+k+\frac{3}{2})}\left(\frac{r_{0}}{2}\right)^{2k}r_{0}^{-\frac{3}{2}},$
$I_{2}(m,r_{0})$ $:=\sum_{k=1}^{\infty}\sum_{k_{1}+k_{2}=k}\left(\frac{\Gamma^{2}(m+\frac{3}{2})}{k_{1}!k_{2}!\Gamma(m+k_{1}+\frac{3}{2})\Gamma(m+k_{2}+\frac{3}{2})}\right)$
$\qquad\times(-1)^{k}\left(\frac{r_{0}}{2}\right)^{2k}\frac{2m+3}{2m+2k+3},$
$I_{3}(m,r_{0})$ $:=\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\frac{\Gamma(m+1)\sqrt{2m+2}}{\Gamma(m+k+1)}\left(\frac{r_{0}}{2}\right)^{2k}r^{-1}_{0},$
$I_{4}(m,r_{0})$ $:=\sum_{k=1}^{\infty}\left(\sum_{k_{1}+k_{2}=k}\frac{\Gamma^{2}(m+1)}{k_{1}!k_{2}!\Gamma(m+k_{1}+1)\Gamma(m+k_{2}+1)}\right)$
$\qquad\times(-1)^{k}\left(\frac{r_{0}}{2}\right)^{2k}\frac{2m+2}{2m+2k+2}.$
Proof.
Set $\nu=m+\frac{1}{2}$, so that $\Gamma(m+k+\frac{3}{2})=\Gamma(\nu+k+1)$ and $\sqrt{2m+3}=\sqrt{2\nu+2}$. Using (2.1), we obtain
$I_{1}=\sqrt{2\nu+2}r_{0}^{-\frac{3}{2}}[\Gamma(\nu+1)(\frac{r_{0}}{2})^{-\nu}J_{\nu}(r_{0})-1].$ (3.4)
Let $\beta=(\frac{r_{0}}{2})^{2}$ and $\gamma_{k}(\nu)=\frac{1}{(\nu+k)(\nu+k-1)\cdots(\nu+1)}$. By the Gamma function recurrence and (2.1),
$|\Gamma(\nu+1)(\frac{r_{0}}{2})^{-\nu}J_{\nu}(r_{0})-1|$ $=|\sum_{k=1}^{\infty}\frac{(-1)^{k}}{k!}\beta^{k}\gamma_{k}(\nu)|$ (3.5)
$\leqslant\exp(\frac{\beta}{\nu})-1\leqslant\frac{\beta}{\nu}\exp(\frac{\beta}{\nu})\leqslant\frac{\beta}{\nu}\exp(\beta),$
where we used $|\gamma_{k}(\nu)|\leqslant\nu^{-k}$. Combining (3.4) and (3.5) yields
$|I_{1}|\leqslant\frac{\sqrt{2\nu+2}}{\nu}r_{0}^{-\frac{3}{2}}\beta\exp(\beta)=\frac{\sqrt{2m+3}}{m+\frac{1}{2}}r_{0}^{-\frac{3}{2}}(\frac{r_{0}}{2})^{2}\exp(\frac{r_{0}}{2})^{2}\rightarrow 0$
as $m\rightarrow\infty$.
For $I_{2}$, rewrite it as
$I_{2}\left(m,r_{0}\right)=(2m+3)\sum_{k=1}^{\infty}\frac{(-1)^{k}\left(r_{0}/2\right)^{2k}}{2m+2k+3}\cdot S_{k}(m),$
where
$S_{k}(m)=\sum_{\begin{subarray}{c}k_{1}+k_{2}=k\\ k_{1},k_{2}\geq 0\end{subarray}}\frac{1}{k_{1}!k_{2}!}\cdot\frac{\Gamma^{2}\left(m+\frac{3}{2}\right)}{\Gamma\left(m+k_{1}+\frac{3}{2}\right)\Gamma\left(m+k_{2}+\frac{3}{2}\right)}.$
For fixed $j\geq 0$, the Gamma recurrence implies
$\frac{\Gamma\left(m+\frac{3}{2}\right)}{\Gamma\left(m+j+\frac{3}{2}\right)}\leq\frac{1}{\left(m+\frac{3}{2}\right)^{j}}.$
Applying this to $S_{k}(m)$ and noting $\sum_{k_{1}+k_{2}=k}\frac{1}{k_{1}!k_{2}!}=\frac{2^{k}}{k!}$, we get
$S_{k}(m)\leq\frac{2^{k}}{k!\left(m+\frac{3}{2}\right)^{k}}.$
Thus,
$\left|I_{2}(m)\right|$ $\leq(2m+3)\sum_{k=1}^{\infty}\frac{1}{2m+2k+3}\cdot\frac{\left(r_{0}^{2}/2\right)^{k}}{k!\left(m+\frac{3}{2}\right)^{k}}$
$\leqslant\sum_{k=1}^{\infty}\frac{\left(r_{0}^{2}/2\right)^{k}}{k!\left(m+\frac{3}{2}\right)^{k}}=\exp(\frac{r_{0}^{2}}{2m+3})-1\rightarrow 0$
as $m\rightarrow\infty$. The same approach applies to $I_{3}$ and $I_{4}$, so their proofs are omitted.
The proof is complete. ∎
Proof of Theorem 3.1.
The proof proceeds in two steps.
Step 1: Construction of the transmission eigenfunction and its properties.
Consider the interior transmission problem (see [5], p. 319):
$\left\{\begin{array}[]{@{}ll@{}}(\Delta+\mathfrak{n}^{2})w=0&\text{in }\displaystyle\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i}),\\ (\Delta+1)v=0&\text{in }\displaystyle\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i}),\\ w=v,\ \dfrac{\partial w}{\partial\nu}=\dfrac{\partial v}{\partial\nu}&\text{on }\partial\left(\displaystyle\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})\right),\end{array}\right$. (3.6)
where $\nu$ denotes the outward unit normal. A pair $(w,v)$ satisfying (3.6) is called a transmission eigenfunction corresponding to the transmission eigenvalue $\omega$. We analyze explicit solutions in both two-dimensional and three-dimensional case, showing that $v(\mathbf{x})$ attains its maximum simultaneously on each $\partial B_{r_{0}}(\mathbf{y}_{i})$ ($i=1,\dots,n$) with arbitrarily large magnitude for large predetermined parameter $m$.
For a fixed positive integer $m$ to be determined, in polar coordinates with $\mathbf{x}\in B_{r_{0}}(\mathbf{y}_{i})$ ($i=1,\dots,n$):
$w(\mathbf{x})=\alpha_{m}^{l}j_{m}(\mathfrak{n}r_{i})Y_{m}^{l}(\theta_{i},\varphi_{i}),\quad v(\mathbf{x})=\beta_{m}^{l}j_{m}(r_{i})Y_{m}^{l}(\theta_{i},\varphi_{i}),\quad-m\leqslant l\leqslant m,$
where $r_{i}=|\mathbf{x}-\mathbf{y}_{i}|$, $\theta_{i}=\arccos((x_{3}-y_{i3})/r_{i})$, $\varphi_{i}=\arctan((x_{2}-y_{i2})/(x_{1}-y_{i1}))$. To satisfy the boundary conditions on $\partial B_{r_{0}}(\mathbf{y}_{i})$, we set $\alpha^{l}_{m}=[j_{m}(r_{0})/j_{m}(\mathfrak{n}r_{0})]\beta^{l}_{m}$ with $\mathfrak{n}$ solving
$\mathfrak{n}j_{m}^{\prime}(\mathfrak{n}r_{0})j_{m}(r_{0})-j_{m}^{\prime}(r_{0})j_{m}(\mathfrak{n}r_{0})=0.$
Normalizing $v$ such that $\|v\|_{L^{2}(B_{r_{0}}(\mathbf{y}_{i}))}=1$ for all $i$, we obtain
$\beta_{m}^{l}=\sqrt{\frac{2}{\pi}}\frac{1}{\sqrt{\int_{0}^{r_{0}}J^{2}_{m+\frac{1}{2}}(r)rdr}}.$
Combining these results yields
$v(\mathbf{x})=\frac{r_{i}^{-\frac{1}{2}}J_{m+\frac{1}{2}}(r_{i})}{\sqrt{\int_{0}^{r_{0}}J_{m+\frac{1}{2}}^{2}(r)rdr}}Y_{m}^{l}(\theta_{i},\varphi_{i}),\quad\mathbf{x}\in B_{r_{0}}(\mathbf{y}_{i}).$ (3.7)
By using recurrence relation for Bessel function, we have
$\left(\frac{J_{m+\frac{1}{2}}(r)}{r^{\frac{1}{2}}}\right)^{\prime}=\frac{mJ_{m+\frac{1}{2}}(r)-rJ_{m+\frac{3}{2}}(r)}{r^{\frac{3}{2}}}.$
Let $m>r_{0}$ and use the fact that $J_{m+\frac{1}{2}}(r)>J_{m+\frac{3}{2}}(r)~,\forall r\in(0,m]$, then $J_{m+\frac{1}{2}}(r)r^{-\frac{1}{2}}$ is increasing on $\left(0,r_{0}\right]$. Consequently, by (2.1), we calculate
$\max_{r_{i}\in\left(0,r_{0}\right]}\frac{{r_{i}^{-\frac{1}{2}}}J_{m+\frac{1}{2}}(r_{i})}{\sqrt{\int_{0}^{r_{0}}J_{m+\frac{1}{2}}^{2}(r)rdr}}=\frac{{r_{0}^{-\frac{1}{2}}}J_{m+\frac{1}{2}}(r_{0})}{\sqrt{\int_{0}^{r_{0}}J_{m+\frac{1}{2}}^{2}(r)rdr}}$ (3.8)
$=\frac{\sum_{k=0}^{\infty}\frac{(-1)^{k}r_{0}^{m+2k}}{k!2^{m+2k+\frac{1}{2}}\Gamma(m+k+\frac{3}{2})}}{\sqrt{\sum_{k=0}^{\infty}\sum_{k_{1}+k_{2}=k}\frac{1}{k_{1}{!}k_{2}{!}\Gamma(m+k_{1}+\frac{3}{2})\Gamma(m+k_{2}+\frac{3}{2})}\frac{(-1)^{k}(r_{0})^{2m+2k+3}}{2^{2m+2k+1}2m+2k+3}}}$
$=\frac{\sqrt{2m+3}r_{0}^{-\frac{3}{2}}+\sum_{k=1}^{\infty}\frac{(-1)^{k}\Gamma(m+\frac{3}{2})\sqrt{2m+3}}{k!\Gamma(m+k+\frac{3}{2})}\left(\frac{r_{0}}{2}\right)^{2k}r_{0}^{-\frac{3}{2}}}{\sqrt{1+\sum_{k=1}^{\infty}\sum_{k_{1}+k_{2}=k}\left(\frac{\Gamma^{2}(m+\frac{3}{2})}{k_{1}!k_{2}!\Gamma(m+k_{1}+\frac{3}{2})\Gamma(m+k_{2}+\frac{3}{2})}\right)\times\frac{(-1)^{k}(r_{0}/2)^{2k}(2m+3)}{(2m+2k+3)}}}$
$=\frac{\sqrt{2m+3}r_{0}^{-\frac{3}{2}}+I_{1}(m,r_{0})}{\sqrt{1+I_{2}(m,r_{0})}}.$
By Lemma 3.5, $I_{1},I_{2}\to 0$ as $m\to\infty$ for fixed $r_{0}$. Thus, there exists $M_{1}(r_{0})$ such that for predetermined parameter $m\geqslant M_{1}$,
$\sqrt{2m+3}r_{0}^{-\frac{3}{2}}\geq\frac{J_{m+\frac{1}{2}}(r_{0})r_{0}^{-\frac{1}{2}}}{\sqrt{\int_{0}^{r_{0}}J_{m+\frac{1}{2}}^{2}(r)rdr}}>\frac{1}{2}\sqrt{2m+3}r_{0}^{-\frac{3}{2}}.$ (3.9)
Combining (2.3), (3.7), (3.9), and spherical harmonic properties from Lemma 2.1, there exist points $\mathbf{y}^{*}_{i}\in\partial B_{r_{0}}(\mathbf{y}_{i})$ satisfying
$|v(\mathbf{y}^{*}_{i})|>\frac{1}{16}\sqrt{2m+3}r_{0}^{-\frac{3}{2}}.$ (3.10)
The two-dimensional case follows analogously. For $\mathbf{x}\in B_{r_{0}}(\mathbf{y}_{i})$ ($i=1,\dots,n$):
$w(\mathbf{x})=\alpha_{m}J_{m}(\mathfrak{n}r_{i})e^{\mathrm{i}m\theta_{i}},\quad v(\mathbf{x})=\beta_{m}J_{m}(r_{i})e^{\mathrm{i}m\theta_{i}},$
with $r_{i}=|\mathbf{x}-\mathbf{y}_{i}|$, $\theta_{i}=\arctan((x_{2}-y_{i2})/(x_{1}-y_{i1}))$. Setting $\alpha_{m}=[J_{m}(r_{0})/J_{m}(\mathfrak{n}r_{0})]\beta_{m}$ and $\beta_{m}=\frac{1}{\sqrt{2\pi}}\frac{1}{\sqrt{\int_{0}^{r_{0}}J^{2}_{m}(r)rdr}}$ yields
$v(\mathbf{x})=\frac{1}{\sqrt{2\pi}}\frac{J_{m}(r_{i})}{\sqrt{\int_{0}^{r_{0}}J_{m}^{2}(r)rdr}}e^{\mathrm{i}m\theta_{i}}.$ (3.11)
For $m>r_{0}$, by (2.1), we obtain
$\max_{r_{i}\in[0,r_{0}]}\frac{J_{m}(r_{i})}{\sqrt{\int_{0}^{r_{0}}J_{m}^{2}(r)rdr}}=\frac{\sqrt{2m+2}r^{-1}_{0}+I_{3}(m,r_{0})}{\sqrt{1+I_{4}(m,r_{0})}}.$ (3.12)
By Lemma 3.5, $I_{3},I_{4}\to 0$ as $m\to\infty$ for fixed $r_{0}$. Then, for any predetermined $m>M_{2}(r_{0})$, there exist points $\mathbf{y}^{*}_{i}\in\partial B{r_{0}}(\mathbf{y}_{i})$ such that
$|v(\mathbf{y}^{*}_{i})|\geqslant\frac{\sqrt{2m+2}r^{-1}_{0}}{2\sqrt{2\pi}}.$ (3.13)
Defining $M=\max\{M_{1}(r_{0}),M_{2}(r_{0}),\lfloor r_{0}\rfloor+1\}$, both (3.10) and (3.13) hold for predetermined parameter $m\geqslant M$.
Step 2: Herglotz approximation and construction of $u_{0}$.
Since $D$ is a bounded Lipschitz domain, it admits arbitrarily close smooth outer approximations; see [3]. In particular, there exists a $C^{\infty}$-smooth bounded domain $D_{r_{0}/4}$ such that
$D\Subset D_{r_{0}/4}\quad\text{and}\quad\sup_{x\in\partial D_{r_{0}/4}}\operatorname{dist}(x,D)<\frac{r_{0}}{4}.$
As established in Theorem 3.1, the sets $D_{r_{0}/4}$ and $B_{r_{0}}(\mathbf{y}_{i})$ are disjoint for all $i$. Then we define $\mathcal{O}=D_{r_{0}/4}\cup\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})$. Consider the interior transmission problem
$\left\{\begin{array}[]{@{}ll@{}}(\Delta+\mathfrak{n}^{2})w_{1}=0&\text{in }\mathcal{O},\\ (\Delta+1)v_{1}=0&\text{in }\mathcal{O},\\ w_{1}=v_{1},\ \dfrac{\partial w_{1}}{\partial\nu}=\dfrac{\partial v_{1}}{\partial\nu}&\text{on }\partial\mathcal{O}.\end{array}\right$. (3.14)
The pair $(w_{1},v_{1})=\chi_{\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})}(w,v)+\chi_{D_{r_{0}/4}}(0,0)$ satisfies (3.14) by construction. Setting $\omega=1$ and applying Lemma 2.2 to $v_{1}=\chi_{\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})}v+\chi_{D_{r_{0}/4}}\cdot 0$, there exists $g\in L^{2}(\mathbb{S}^{d-1})$ such that the Herglotz function
$H_{g}(\mathbf{x})=\int_{\mathbb{S}^{d-1}}g(\theta)\exp(\mathrm{i}\mathbf{x}\cdot\theta)d\theta$ (3.15)
satisfies
$\left\|H_{g}-\left(\chi_{\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})}v+\chi_{D_{r_{0}/2}}\cdot 0\right)\right\|_{H^{5}(\mathcal{O})}<\varepsilon.$ (3.16)
This implies
$\|H_{g}-v\|_{H^{5}(B_{r_{0}}(\mathbf{y}_{i}))}<\varepsilon\quad(i=1,\dots,n)\quad\text{and}\quad\|H_{g}\|_{H^{5}(D)}<\varepsilon.$ (3.17)
Define
$u_{0}(\mathbf{x},t)=H_{g}(\mathbf{x})\exp(-\mathrm{i}t).$ (3.18)
Since $H_{g}(\mathrm{x})$ satisfies $(\Delta+1)H_{g}(\mathrm{x})=0$ in $\mathbb{R}^{d}$, a straightforward calculation shows that the function $u_{0}$ defined above satisfies
$\mathrm{i}\partial_{t}u_{0}+\Delta u_{0}=0\quad\text{in }\mathbb{R}^{d}\times[0,+\infty),$
which implies that Properties 1 and 2 hold. Furthermore, by the Sobolev embedding theorem and the estimate (3.17), there exist constants $C_{1}$ (depending on $D$ and $d$) and $C_{2}$ (depending on $r_{0}$, $D$, and $d$) such that
$\|H_{g}\|_{C^{1,\frac{1}{2}}(\overline{D})}\leqslant C_{1}\varepsilon,$ (3.19)
and
$\|H_{g}-v\|_{L^{\infty}(B_{r_{0}}(\mathbf{y}_{i}))}\leqslant C_{2}\|H_{g}-v\|_{H^{5}(B_{r_{0}}(\mathbf{y}_{i}))}\leqslant C_{2}\varepsilon\quad(i=1,\dots,n).$ (3.20)
Combining (3.10), (3.13), (3.18), (3.19), and (3.20), together with the continuity of $H_{g}$ and $v$, we conclude that Properties 3 and 4 are verified.
The proof is complete. ∎
Remark 3.6.
The approximation of $\chi_{\bigcup_{i=1}^{n}B_{r_{0}}(\mathbf{y}_{i})}v+\chi_{D_{r_{0}/2}}\cdot 0$ in the $H^{5}$ norm required in (3.16) is valid due to the $C^{\infty}$ regularity of $\partial D_{r_{0}/2}$. This construction serves a critical purpose: the resulting estimate will be employed in the subsequent section to separately control the source term $F_{1}$ in equation (4.1), the source term $F_{2}$ in equation (5.1), and all spatial derivatives $\partial_{\mathbf{x}}^{\alpha}u_{0}$ for $|\alpha|\leqslant 3$. Detailed justification for this approach is provided in (4.4) and Lemma 5.1.
The initial and boundary data for problem (1.1) are formally specified as:
$\phi(\mathbf{x})=\left.u_{0}(\mathbf{x},t)\right|_{\Omega\times\{t=0\}}=H_{g}(\mathbf{x})|_{\Omega},$ (3.21)
$\psi(\mathbf{x},t)=u_{0}\left.(\mathbf{x},t)\right|_{\partial\Omega\times[0,T]}=\exp(-\mathrm{i}t)H_{g}(\mathbf{x})|_{\partial\Omega\times[0,T]}.$
Similarly, the initial data for problem (1.2) are defined as:
$\Phi(\mathbf{x})=\left.u_{0}(\mathbf{x},t)\right|_{\mathbb{R}^{d}\times\{t=0\}}=H_{g}(\mathbf{x})|_{\mathbb{R}^{d}},$ (3.22)
Remark 3.7.
By Remark (2.3) and the assumption that $\partial\Omega$ is Lipschitz, we have $\phi\in C^{\infty}(\Omega)$, $\psi\in C^{\infty}([0,T];C(\partial\Omega))$, and $\Phi\in C^{\infty}(\mathbb{R}^{d})$.
Remark 3.8.
It is essential to recognize that the constructions in (3.21) and (3.22) remain abstract at this stage. The function $u_{0}$–and consequently all associated data–depends critically on the yet-to-be-determined parameters $\varepsilon$, $m$, $r_{0}$, and the configuration of points $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ on $\partial D$. A concrete realization of $u_{0}$ requires explicitly specifying these parameters, which are intrinsically determined by the prescribed large constant $\mathcal{M}$ and the spatial locations of the points $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$. In the next two sections, we will strategically choose $\varepsilon$, $m$, and $r_{0}$ in terms of $\mathcal{M}$ and the given boundary point configuration, ultimately yielding a well-defined $u_{0}$ and corresponding initial and boundary inputs suitable for rigorous analysis.
4. Proof of Theorem 1.1
In this section, we present the complete proof of Theorem 1.1. Our strategy involves analyzing an auxiliary problem (4.1) with zero initial-boundary conditions and a source term, derived from the linear Schrödinger equation (1.1). We establish well-posedness for this auxiliary problem via the standard Galerkin method and demonstrate that its solution exhibits small Hölder norm on $\Omega\setminus\overline{D}$. Finally, by integrating Theorems 3.1 and 4.1 and a delicate balancing of the parameters $\varepsilon$, $r_{0}$, and $m$, we precisely generate the localized gradient amplification behavior central to our main result.
Let $\mathcal{U}:\Omega\times[0,T]\rightarrow\mathbb{C}$ satisfy
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}+L_{1}\mathcal{U}=F_{1}&\text{in }\Omega\times(0,T],\\ \mathcal{U}(\mathbf{x},0)=0&\text{in }\Omega,\\ \mathcal{U}(\mathbf{x},t)=0&\text{on }\partial\Omega\times[0,T],\end{cases}$ (4.1)
where the source term
$F_{1}(\mathbf{x},t)=\nabla\cdot\left((\mathbf{I}-\mathbf{A}_{1})\nabla u_{0}\right).$ (4.2)
Recalling that $\mathbf{A}_{1}-\mathbf{I}$ has compact support in $D$, we obtain
$\operatorname{supp}F_{1}(\cdot,t)\subset D,\quad\forall t\in[0,T].$
Furthermore, combining (3.17), (3.18), (4.2), and the assumption $\mathbf{A}_{1}\in W^{2,\infty}(\Omega)$, we obtain the following estimate for $F_{1}$:
$\|F_{1}\|_{L^{\infty}(0,T;H^{1}(\Omega))}\leqslant C\varepsilon,$ (4.3)
and
$\|F_{1}\|_{H^{2}(0,T;L^{2}(\Omega))}+\|F_{1}\|_{H^{1}(0,T;H^{1}(\Omega))}\leqslant C\|H_{g}\|_{H^{3}(D)}\leqslant C\varepsilon.$ (4.4)
Here, the spatial derivatives of $u_{0}$ appearing in $F_{1}$ are of order at most two, whereas the space–time Sobolev norm estimate requires third-order spatial derivatives of $H_{g}$ due to the presence of spatial gradients in the $H^{1}(\Omega)$ norm. Hence, the $H^{3}$ norm of $H_{g}$ suffices to control all terms in the estimate. The constant $C$ depends on $d$, $T$, $D$, and $\|\mathbf{A}_{1}\|_{W^{2,\infty}(\Omega)}$.
Next, we obtain the well-posedness for auxiliary problem (4.1) by energy estimates and Galerkin method.
Theorem 4.1.
For any $T$ with $0<T<\infty$, problem (4.1) admits a unique solution $\mathcal{U}$ satisfying
$\mathcal{U}\in L^{\infty}\left(0,T;H^{3}_{\mathrm{loc}}(\Omega)\right)$
with the following properties:
  1. (1)
    Interior improved Hölder regularity on $\Omega:$ for any Lipschtiz domain $\Omega^{\prime}\Subset\Omega$,
    $\mathcal{U}\in L^{\infty}(0,T;C^{1,\frac{1}{2}}(\overline{\Omega}^{\prime})).$
  2. (2)
    Smallness estimate:
    $\|\mathcal{U}\|_{L^{\infty}(0,T;C^{1,\frac{1}{2}}(\overline{\Omega}^{\prime}))}\leqslant C_{3}\varepsilon,$
    where the constant $C_{3}$ depends on where the constant $C_{3}$ depends on $d$, $T$, $\theta$, $D$, $\Omega^{\prime}$, $\Omega$, $\|\mathbf{A}_{1}\|_{W^{2,\infty}(\Omega)}$$\|\mathbf{A}_{1}\|_{C^{1,1}(\overline{\Omega})}$.
Proof.
Step 1: a priori estimates for problem (4.1).
Introduce the notation
$A[\mathcal{U},\mathcal{V};t]:=\int_{\Omega}\sum_{i,j=1}^{d}a_{1}^{ij}(\cdot)\,\mathcal{U}_{x_{i}}\,\overline{\mathcal{V}_{x_{j}}}.$
Multiplying (4.1) by $\overline{\mathcal{U}}$, using integration by parts, we obtain
$\mathrm{i}\int_{\Omega}\partial_{t}\mathcal{U}\,\overline{\mathcal{U}}\,dx-\int_{\Omega}a^{ij}_{1}(x)\partial_{i}\mathcal{U}\,\overline{\partial_{j}\mathcal{U}}\,dx=\int_{\Omega}F_{1}\,\overline{\mathcal{U}}\,dx.$ (4.5)
Taking the complex conjugate of (4.5), subtracting it from (4.5), taking the imaginary part, and applying the Cauchy–Schwarz inequality yields:
$\frac{d}{dt}\|\mathcal{U}(t)\|_{L^{2}(\Omega)}^{2}\leq C\Bigl(\|\mathcal{U}(t)\|_{L^{2}}^{2}+\|F_{1}(t)\|_{L^{2}}^{2}\Bigr).$
By Grönwall’s inequality,
$\|\mathcal{U}(t)\|_{L^{2}(\Omega)}^{2}\leq C\int_{0}^{t}\|F_{1}(s)\|_{L^{2}(\Omega)}^{2}\,ds,$ (4.6)
for all $0\leq t\leq T$.
Multiplying (4.1) by $\overline{\partial_{t}\mathcal{U}}$, and integrating by parts, we obtain
$\mathrm{i}\int_{\Omega}|\partial_{t}\mathcal{U}|^{2}\,dx-\int_{\Omega}a^{ij}_{1}(x)\partial_{i}\mathcal{U}\,\partial_{t}\overline{\partial_{j}\mathcal{U}}\,dx=\int_{\Omega}F_{1}\,\overline{\partial_{t}\mathcal{U}}\,dx.$ (4.7)
Note that
$\int_{\Omega}\overline{a^{ij}_{1}(x)\partial_{i}\mathcal{U}\,\partial_{t}\overline{\partial_{j}\mathcal{U}}}\,dx=\int_{\Omega}a^{ji}_{1}(x)\,\overline{\partial_{i}\mathcal{U}}\,\partial_{t}\partial_{j}\mathcal{U}\,dx=\int_{\Omega}a^{ij}_{1}(x)\,\partial_{t}\partial_{i}\mathcal{U}\,\overline{\partial_{j}\mathcal{U}}\,dx.$
Taking the complex conjugate of (4.7), adding it back, and taking the real part gives
$\frac{d}{dt}A[\mathcal{U},\mathcal{U};t]=-2\Re\int_{\Omega}F_{1}(t)\,\overline{\partial_{t}\mathcal{U}(t)}\,dx.$
Integrating over $[0,t]$, using integration by parts in time, the identity $\mathcal{U}(0)=0$, and (4.6), we calculate
$A[\mathcal{U},\mathcal{U};t]$ $=A[\mathcal{U},\mathcal{U};0]-2\Re\int_{0}^{t}\int_{\Omega}F_{1}(x,s)\,\overline{\partial_{s}\mathcal{U}(x,s)}\,dx\,ds$
$=-2\Re\int_{\Omega}F_{1}(x,t)\overline{\mathcal{U}(x,t)}\,dx+2\Re\int_{0}^{t}\int_{\Omega}\partial_{s}F_{1}(x,s)\,\overline{\mathcal{U}(x,s)}\,dx\,ds$
$\leq C\left(\|F_{1}(t)\|_{L^{2}}^{2}+\|\mathcal{U}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|F_{1}^{\prime}(s)\|_{L^{2}}^{2}+\|\mathcal{U}(s)\|_{L^{2}}^{2}\,ds\right)$
$\leq C\|F_{1}\|_{W^{1,2}(0,T;L^{2}(\Omega))}^{2}.$
Thus,
$\|\mathcal{U}(t)\|_{H^{1}_{0}(\Omega)}\leq C\,\|F_{1}\|_{W^{1,2}(0,T;L^{2}(\Omega))},$ (4.8)
for all $t\in[0,T]$.
Next, set $\mathcal{U}_{1}:=\partial_{t}\mathcal{U}$. Then $\mathcal{U}_{1}$ satisfies
$\begin{cases}\mathrm{i}\partial_{t}\mathcal{U}_{1}+L_{1}\mathcal{U}_{1}=F_{1}^{\prime},&\text{in }\Omega\times(0,T],\\[5.69054pt] \mathcal{U}_{1}(x,0)=-\mathrm{i}F_{1}(x,0),&\text{in }\Omega,\\[5.69054pt] \mathcal{U}_{1}(x,t)=0,&\text{on }\partial\Omega\times[0,T].\end{cases}$ (4.9)
Since $F_{1}$ has compact support, $\mathcal{U}_{1}(x,0)|_{\partial\Omega}=-\mathrm{i}F_{1}(x,0)|_{\partial\Omega}=0$, which is compatible with the boundary condition. Repeating the same procedure as above, multiplying (4.9) by $\overline{\mathcal{U}_{1}}$, integrating by parts, we obtain
$\mathrm{i}\int_{\Omega}\partial_{t}\mathcal{U}_{1}\,\overline{\mathcal{U}_{1}}\,dx-\int_{\Omega}a^{ij}_{1}(x)\partial_{i}\mathcal{U}_{1}\,\overline{\partial_{j}\mathcal{U}_{1}}\,dx=\int_{\Omega}F_{1}^{\prime}\,\overline{\mathcal{U}_{1}}\,dx.$ (4.10)
Taking the conjugate of (4.10), subtracting, and taking imaginary parts again yields
$\frac{d}{dt}\|\mathcal{U}_{1}(t)\|_{L^{2}}^{2}\leq C\Bigl(\|\mathcal{U}_{1}(t)\|_{L^{2}}^{2}+\|F_{1}^{\prime}(t)\|_{L^{2}}^{2}\Bigr),$
and by Grönwall’s inequality,
$\|\mathcal{U}_{1}(t)\|_{L^{2}}^{2}\leq C\int_{0}^{t}\|F_{1}^{\prime}(s)\|_{L^{2}}^{2}\,ds,$ (4.11)
for all $0\leq t\leq T$. Multiplying (4.9) by $\overline{\partial_{t}\mathcal{U}_{1}}$, integrating by parts, we obtain
$\mathrm{i}\int_{\Omega}|\partial_{t}\mathcal{U}_{1}|^{2}\,dx-\int_{\Omega}a^{ij}_{1}(x)\partial_{i}\mathcal{U}_{1}\,\partial_{t}\overline{\partial_{j}\mathcal{U}_{1}}\,dx=\int_{\Omega}F_{1}^{\prime}\,\overline{\partial_{t}\mathcal{U}_{1}}\,dx.$
Proceeding as before, we derive
$\frac{d}{dt}A[\mathcal{U}_{1},\mathcal{U}_{1};t]=-2\Re\int_{\Omega}F_{1}^{\prime}(t)\,\overline{\partial_{t}\mathcal{U}_{1}(t)}\,dx.$
Integrating over $[0,t]$, integrating by parts in $t$, and using $\mathcal{U}_{1}(0)=-\mathrm{i}F_{1}(x,0)$ together with (4.11), we obtain
$A[\mathcal{U}_{1},\mathcal{U}_{1};t]$ $=A[\mathcal{U}_{1},\mathcal{U}_{1};0]-2\Re\int_{0}^{t}\int_{\Omega}F_{1}^{\prime}(x,s)\,\overline{\partial_{s}\mathcal{U}_{1}(x,s)}\,dx\,ds$ (4.12)
$=A[\mathcal{U}_{1},\mathcal{U}_{1};0]-2\Re\int_{\Omega}F_{1}^{\prime}(x,t)\overline{\mathcal{U}_{1}(x,t)}-F_{1}^{\prime}(x,0)\overline{\mathcal{U}_{1}(x,0)}\,dx$
$\quad+2\Re\int_{0}^{t}\int_{\Omega}F_{1}^{\prime\prime}(x,s)\,\overline{\mathcal{U}_{1}(x,s)}\,dx\,ds$
$\leq C\Bigl(\|\nabla F_{1}(0)\|_{L^{2}}^{2}+\|F_{1}^{\prime}(t)\|_{L^{2}}^{2}+\|F_{1}^{\prime}(0)\|_{L^{2}}^{2}+\|\mathcal{U}_{1}(t)\|_{L^{2}}^{2}+\|\mathcal{U}_{1}(0)\|_{L^{2}}^{2}$
$\quad+\int_{0}^{t}\|F_{1}^{\prime\prime}(s)\|_{L^{2}}^{2}+\|\mathcal{U}_{1}(s)\|_{L^{2}}^{2}\,ds\Bigr)$
$\leq C\Bigl(\|F_{1}\|_{H^{1}(0,T;H^{1}(\Omega))}^{2}+\|F_{1}\|_{H^{2}(0,T;L^{2}(\Omega))}^{2}\Bigr).$
Combining (4.11) and (4.12), and using the coercivity $A[\mathcal{U}_{1},\mathcal{U}_{1};t]\geq\theta\|\nabla\mathcal{U}_{1}(t)\|_{L^{2}}^{2}$, we deduce
$\|\mathcal{U}_{1}(t)\|_{H^{1}(\Omega)}^{2}\leq C\Bigl(\|F_{1}\|_{H^{1}(0,T;H^{1}(\Omega))}^{2}+\|F_{1}\|_{H^{2}(0,T;L^{2}(\Omega))}^{2}\Bigr),$ (4.13)
for all $0\leq t\leq T$.
Step 2: existence, uniqueness and Hölder regularity for problem (4.1).
Using the Galerkin method and the a priori estimate (4.8), we obtain a weak solution
$\mathcal{U}\in L^{\infty}(0,T;H^{1}_{0}(\Omega))$
to (4.1). Uniqueness follows from Grönwall’s inequality. By (4.4) and (4.13), we also have
$\partial_{t}\mathcal{U}\in L^{\infty}(0,T;H_{0}^{1}(\Omega))$
and
$\|\partial_{t}\mathcal{U}\|_{L^{\infty}(0,T;H_{0}^{1}(\Omega))}\leq C\left(\|F_{1}\|_{H^{1}(0,T;H^{1}(\Omega))}+\|F_{1}\|_{H^{2}(0,T;L^{2}(\Omega))}\right)\leq C\varepsilon.$
Then, by the above estimate and (4.3), the solution $\mathcal{U}(t)\in H^{1}(\Omega)$ satisfies
$L_{1}\mathcal{U}(t)=F_{1}(t)-\mathrm{i}\,\partial_{t}\mathcal{U}(t)\in H^{1}(\Omega)$
in the weak sense for almost every $t\in[0,T]$. Since $\mathbf{A}_{1}\in C^{1,1}(\overline{\Omega})$, interior elliptic regularity (Theorem 8.10 in [9]) implies that, for any $\Omega^{\prime}\Subset\Omega\setminus\overline{D}$,
$\mathcal{U}(t)\in H^{3}(\Omega^{\prime}),\qquad\|\mathcal{U}(t)\|_{H^{3}(\Omega^{\prime})}\leq C\Bigl(\|\partial_{t}\mathcal{U}(t)\|_{H^{1}(\Omega)}+\|F_{1}(t)\|_{H^{1}(\Omega)}\Bigr)\leq C\varepsilon,$
for almost every $t\in[0,T]$. Taking the supremum over $t\in[0,T]$ yields
$\|\mathcal{U}\|_{L^{\infty}(0,T;H^{3}(\Omega^{\prime}))}\leq C\varepsilon.$
Finally, by the Sobolev embedding theorem, properties (1) and (2) follow.
The proof is complete. ∎
We now proceed to establish Theorem 1.1.
Proof of Theorem 1.1..
In the three-dimensional case, according to Remark 3.8, we select predetermined parameters $r_{0}$, $\varepsilon$, and $m$ as follows:
$\hfill r_{0}=\min\left\{\,\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|,\ \frac{1}{3}\operatorname{dist}(D,\partial\Omega)\,\right\},$ (4.14)
$\hfill\varepsilon=\min\left\{\frac{1}{C_{1}+C_{2}(r_{0})+C_{3}},1\right\},$
$\hfill m=\max\left\{\left\lfloor 512r_{0}^{3}\right\rfloor+1,M(r_{0})\right\},$
where the constants $C_{1}$, $C_{2}(r_{0})$, and $M(r_{0})$ are specified in Theorem 3.1, and $C_{3}$ is given by Theorem 4.1. Then, we choose the initial data $\phi$ and boundary data $\psi$ for problem (1.1) as specified in (3.21), where the corresponding $u_{0}$ is determined by the parameters $r_{0}$, $\varepsilon$, and $m$ in (4.14). Let $\mathcal{U}$ denote the unique solution to auxiliary problem (4.1). A direct calculation verifies that
$u:=u_{0}+\mathcal{U}$ (4.15)
constitutes the unique solution to problem (1.1). By properties 2 of Theorem 4.1 and the smoothness of $u_{0}$, it follows that $u\in L^{\infty}(0,T;C^{1,\frac{1}{2}}(\overline{\Omega^{\prime}}))$ for any $\Omega^{\prime}\Subset\Omega$ with Lipschitz boundary. Applying the mean value theorem to (3.1) and (3.3) in Theorem 3.1 yields points $\D¯\mathbf{z}_{i}\in B_{3r_{0}}(\mathbf{x}_{i})\backslash\overline${D} ($i=1,2,\dots,n$) such that for all $t\in[0,T]$:
$|\nabla u_{0}(\mathbf{z}_{i},t)|\geqslant\frac{\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}-\left(C_{1}+C_{2}(r_{0})\right)\varepsilon}{3r_{0}}.$ (4.16)
Combining (4.14), (4.15), (4.16), and properties (2) of Theorem 4.1, we obtain
$|\nabla u(\mathbf{z}_{i},t)|$ $\geqslant\frac{\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}-\left(C_{1}+C_{2}(r_{0})\right)\varepsilon}{3r_{0}}-C_{3}\varepsilon$ (4.17)
$>\frac{\tfrac{1}{16}\sqrt{1024r_{0}^{3}}\,r_{0}^{-3/2}-1}{3r_{0}}-1=\frac{1}{3r_{0}}-1>\mathcal{M},$
for a.e. $t\in[0,T]$. Here we choose $m\geqslant\left\lfloor 512r_{0}^{3}\right\rfloor+1$ in (4.14) to guarantee $\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}\geqslant 2$. On the other hand, using prpoerties (3) in Theorem 3.1 and above estimate, we calculate
[…]
By (4.17) and (4.18), properties (1) and (2) in Theorem 1.1 are verified. Furthermore, under the above construction, the localized gradient amplification points are confined within the balls $B_{3r_{0}}(\mathbf{x}_{i}),i=1,2,...,n,$ which implies that
[…]
as $\mathcal{M}\to\infty$, for a.e. $t\in[0,T]$. Properties (3) in Theorem 1.1 is verified. The two-dimensional case follows analogously.
The proof is complete. ∎
5. Proof of Theorem 1.2
In this section, we apply identical methodology to the nonlinear Schrödinger equation (1.2), proving Theorem 1.2. Building upon the linear framework, we consider a nonlinear auxiliary problem (5.1) with zero initial conditions, derived from the problem (1.2). Using a fixed-point argument in a suitably chosen Banach space, we establish existence and uniqueness of equation (5.1) while demonstrating the solution’s Hölder norm remains small. Strategic parameter tuning of $\varepsilon$, $r_{0}$, and $m$ completes the proof.
Let $\mathscr{U}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{C}$ satisfy
[…]
where the following conditions hold:
  1. (1)
    $\hat{L}_{2}\mathscr{U}:=\nabla\cdot(\mathbf{A}_{2}\nabla\mathscr{U})+\hat{c}\,\mathscr{U}.$
  2. (2)
    $\hat{c}(\mathbf{x},t)=c(\mathbf{x},t)+\sum_{k=2}^{l_{0}}k\,\alpha_{k}(\mathbf{x},t)\,u_{0}^{\,k-1}(\mathbf{x},t).$
  3. (3)
    The nonlinear term is given by
    $\hat{N}(\mathscr{U})=\sum_{k=2}^{l_{0}}\left(\alpha_{k}\sum_{i=2}^{k}\binom{k}{i}u_{0}^{\,k-i}\mathscr{U}^{i}\right).$ (5.2)
  4. (4)
    The source term is given by
    $F_{2}(\mathbf{x},t)=\nabla\cdot\big(\mathbf{I}-\mathbf{A}_{2}\big)\nabla u_{0}-c\,u_{0}-\sum_{k=2}^{l_{0}}\alpha_{k}u_{0}^{k}.$ (5.3)
Next we give some properties for the coefficients, nonlinear term and source term in (5.1) which will be used in the proof of Theorem 5.2.
Lemma 5.1.
The coefficients, nonlinear term and source term in equation (5.1) satisfy the following properties:
  1. (1)
    $\operatorname{supp}\hat{c}(\cdot,t)$, $\operatorname{supp}\hat{N}(\mathscr{U})(\cdot,t)$ and $\operatorname{supp}F_{2}(\cdot,t)\subset D$ for all $t\in[0,T]$.
  2. (2)
    For a.e. $t\in[0,T]$,
    $\|\hat{c}(\cdot,t)\|_{W^{2,\infty}(D)}\leq C,\qquad\|F_{2}(\cdot,t)\|_{H^{3}(D)}\leq C\,\varepsilon,$
    where the constant $C$ depends only on $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$.
Proof.
Properties 1 follows immediately from the compact support in $D$ of $\mathbf{I}-\mathbf{A}_{2}(\cdot,t)$, $c(\cdot,t)$, $\alpha_{k}(\cdot,t)$ for $t\in[0,T]$, $k=2,\dots,l_{0}$ combined with the expressions defining $\hat{c}$, $\hat{N}(\mathscr{U})$ and $F_{2}$.
For properties 2, note that $\alpha_{k}$ are smooth, so all their derivatives are bounded in $D$. It remains to show $\partial_{\mathbf{x}}^{\alpha}u_{0}\in L^{\infty}(D)$ for all multi-indices $\alpha$ with $|\alpha|\leqslant 3$. By the Sobolev embedding theorem, (3.17), and (3.18), we have
$\|\partial_{\mathbf{x}}^{\alpha}u_{0}\|_{L^{\infty}(D)}\leqslant C\|\partial_{\mathbf{x}}^{\alpha}u_{0}\|_{H^{2}(D)}\leqslant C\|H_{g}\|_{H^{5}(D)}\leqslant C\varepsilon\leqslant C,\quad\forall|\alpha|\leqslant 3,\quad\forall t\in[0,T],$ (5.4)
where the final inequality holds since $\varepsilon\leqslant 1$ (See Theorem 3.1 and Remark 3.2). Consequently,
$\|\hat{c}(\cdot,t)\|_{W^{2,\infty}(D)}\leqslant C\left(\|c_{2}(\cdot,t)\|_{W^{2,\infty}(D)}+\sum_{k=2}^{l_{0}}\|u_{0}\|_{W^{2,\infty}(D)}^{k-1}\right)\leqslant C\left(1+\sum_{k=1}^{l_{0}-1}\varepsilon^{k}\right)\leqslant C,$
for a.e. $t\in[0,T]$.
For the source term, using the Banach algebra property of $H^{3}(D)$ and (3.17), (5.3), and since $\varepsilon\leq 1$, we obtain
$\|F_{2}(\cdot,t)\|_{H^{3}(D)}\leq C\sum_{k=1}^{l_{0}}\|H_{g}\|_{H^{5}(D)}^{k}\leq C\sum_{k=1}^{l_{0}}\varepsilon^{k}\leq C\varepsilon,$ (5.5)
for a.e. $t\in[0,T]$. ∎
We now establish the well-posedness of problem (5.1) via Banach’s fixed point theorem, combined with an estimate for the nonlinear term $F_{2}$. This approach yields an explicit characterization of the solution’s existence time $T$ in terms of the parameter $\varepsilon$. We show that $T$ admits a uniform positive lower bound independent of $\varepsilon$ and, moreover, becomes arbitrarily large for sufficiently small $\varepsilon$. Furthermore, we prove that the Hölder norm of the solution is bounded by $C\sqrt{\varepsilon}$ for some constant $C>0$.
Theorem 5.2.
There exists a time $T>0$ ((given below by (5.21))), depending on $\varepsilon$, $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, such that problem (5.1) admits a unique solution $\mathscr{U}$ satisfying:
$\mathscr{U}\in L^{\infty}(0,T;H^{3}(\mathbb{R}^{d}))$
with the following properties:
  1. (1)
    Uniform bounds for $T$: TT_lower , where $T_{\mathrm{lower}}$ is given by (5.22) and independent of $\varepsilon$.
  2. (2)
    Hölder regularity: U L^([0,T]; C^1,12(R^d)).
  3. (3)
    Smallness estimate: U_L^(0,T;C^1,12(R^d)) C_6 ε, where the constant $C_{6}$ depends only on $d$.
Proof.
The proof proceeds in four steps.
Step 1: Setup of fixed-point framework.
Define the Banach space
$X:=\left\{{u}\in L^{\infty}\left(0,T;H^{1}\left(\mathbb{R}^{d}\right)\right)\right\}$
with norm
$\|{u}\|:=\underset{0\leq t\leq T}{\operatorname{ess}\sup}\|{u}(t)\|_{H^{1}\left(\mathbb{R}^{d}\right)}.$
Introduce the stronger norm
$|{u}|||:=\underset{0\leq t\leq T}{\operatorname{ess}\sup}\|{u}(t)\|_{H^{3}\left(\mathbb{R}^{d}\right)},$
and the closed subset
[…]
For $\mathscr{V}\in X_{\varepsilon}$, define $\mathscr{U}=\mathscr{F}[\mathscr{V}]$ as the solution to the linear initial-value problem
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{U}+\hat{L}_{2}\mathscr{U}=F_{2}-\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.6)
Step 2: A priori estimate for the problem (5.6).
For problem (5.6), we claim that we have the estimate
$|\mathscr{U}|||\leqslant C_{4}(T)T\varepsilon,$ (5.7)
where constant $C_{4}(T)$ depends continuously on $T$, $d$, $l_{0}$, $D$, $\|\mathbf{A}_{2}\|_{L^{\infty}(0,T)}$, $\|c\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, and $\|\alpha_{k}\|_{L^{\infty}(0,T;W^{3,\infty}(D))}$, but not on $\varepsilon$. Moreover, $C_{4}(T)$ is monotonically increasing in $T$ with $C_{4}(0)<\infty$.
For any multi-index $\beta$ with $0\leq|\beta|\leq 3$, apply $\partial_{\mathbf{x}}^{\beta}$ to (5.6) to obtain
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+\hat{L}_{2}\partial^{\beta}_{\mathrm{x}}\mathscr{U}+[\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]\mathscr{U}=\partial^{\beta}_{\mathrm{x}}F_{2}-\partial^{\beta}_{\mathrm{x}}\hat{N}(\mathscr{V})&&\text{on }\mathbb{R}^{d}\times(0,T],\\ &\partial^{\beta}_{\mathrm{x}}\mathscr{U}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d},\end{aligned}\right$. (5.8)
where $\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]=\partial_{\mathbf{x}}^{\beta}\hat{L}_{2}-\hat{L}_{2}\partial_{\mathbf{x}}^{\beta}$ denotes the commutator, which is a differential operator of order at most 2 in spatial derivatives. Multiply the first line in (5.8) with $\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}$, use integration by parts one obtains,
$\mathrm{i}\int_{\mathbb{R}^{d}}\partial_{t}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}-\int_{\mathbb{R}^{d}}a^{ij}_{2}(t)\partial_{i}\partial^{\beta}_{\mathrm{x}}\mathscr{U}\overline{\partial_{j}\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ $+\int_{\mathbb{R}^{d}}[\partial_{\mathbf{x}}^{\beta},\hat{L}_{2}]\mathscr{U}\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}$ (5.9)
$=\int_{\mathbb{R}^{d}}(\partial^{\beta}_{\mathrm{x}}F_{2}-\partial^{\beta}_{\mathrm{x}}\hat{N}(\mathscr{V}))\overline{\partial^{\beta}_{\mathrm{x}}\mathscr{U}}d\mathbf{x}.$
Then taking conjugate in (5.9), subtracting, taking the imaginary part, using integration by parts and Cauchy Schwartz inequality yields
[…]
Summing over $\beta$ with $0\leqslant|\beta|\leqslant 3$ and applying Gronwall’s inequality, we obtain that
[…]
for a.e. $0\leqslant t\leqslant T$.
Next, we estimate $\hat{N}(\mathscr{V})||_{H^{3}(\mathbb{R}^{d})}$. By (5.2), $\operatorname{supp}\hat{N}(\mathscr{V}(t))\subset D$. Since $\mathscr{V}\in X_{\varepsilon}$, then $\mathscr{V}(t)||_{H^{3}(\mathbb{R}^{d})}\leqslant\sqrt{\varepsilon}$ for a.e. $t\in[0,T]$. By (3.17) and (3.18), we have $u_{0}(t)||_{H^{3}(D)}\leqslant C\varepsilon$ for all $t\in[0,T]$. Combine the above results and $\varepsilon<1$, we calculate
[…]
$\leqslant C\sum_{k=2}^{l_{0}}\sum_{i=2}^{k}\varepsilon^{\frac{i}{2}}\varepsilon^{k-i}\leqslant C\varepsilon,$
for a.e. $t\in[0,T]$. Combining estimates (5.5), (5.11), and (5.12), we obtain the desired bound (5.7). By the Hahn–Banach theorem, together with the preceding estimates (see, for example, Theorem 3.2 in [18]), we conclude that the linear problem (5.6) admits a unique solution $\mathscr{U}\in L^{\infty}(0,T;H^{3}(\mathbb{R}^{d})).$
Step 3: Contraction mapping argument and well-posedness of problem (5.1).
By properly choosing $T$, we show that
$\mathscr{F}:X_{\varepsilon}\to X_{\varepsilon}$ (5.13)
and
$\|\mathscr{F}[\mathscr{V}_{1}]-\mathscr{F}[\mathscr{V}_{2}]\|\leqslant\frac{1}{2}\|\mathscr{V}_{1}-\mathscr{V}_{2}\|,\ \forall\mathscr{V}_{1},\mathscr{V}_{2}\in X_{\varepsilon}.$ (5.14)
To ensure that (5.13) holds, we select $T$ such that, according to (5.7),
[…]
This condition is equivalent to
$C_{4}(T)T\leqslant\frac{1}{\sqrt{\varepsilon}}.$ (5.15)
For (5.14), let $\mathscr{U}_{1}=\mathscr{F}[\mathscr{V}_{1}]$, $\mathscr{U}_{2}=\mathscr{F}[\mathscr{V}_{2}]$, and $\mathscr{W}=\mathscr{U}_{1}-\mathscr{U}_{2}$. Then $\mathscr{W}$ satisfies
$\left\{\begin{aligned} &\mathrm{i}\partial_{t}\mathscr{W}+\hat{L}_{2}\mathscr{W}=-\big(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2})\big)&&\text{in }\mathbb{R}^{d}\times(0,T],\\ &\mathscr{W}(\mathbf{x},0)=0&&\text{in }\mathbb{R}^{d}.\end{aligned}\right$. (5.16)
Repeating the procedure to obtain estimate (5.10), for problem (5.16), we have, for any multiindex $\gamma$ with $0\leqslant|\gamma|\leqslant 1$,
[…]
where $\partial_{\mathbf{x}}^{\gamma},\hat{L}_{2}]=\partial_{\mathbf{x}}^{\gamma}\hat{L}_{2}-\hat{L}_{2}\partial_{\mathbf{x}}^{\gamma}$ denotes the commutator, which is a differential operator of order 0 in spatial derivatives. By using (5.2), the fact that $\mathscr{V}_{1},\mathscr{V}_{2}\in X_{\varepsilon}$ implies $\|\mathscr{V}_{i}\|_{W^{1,\infty}(\mathbb{R}^{d})}\leqslant C\sqrt{\varepsilon}$ for $i=1,2$ and Lemma 5.1, we calculate the term $\partial^{\gamma}_{\mathrm{x}}(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2}))||^{2}_{L^{2}(\mathbb{R}^{d})}$ as follows,
$\quad\int_{\mathbb{R}^{d}}|\partial^{\gamma}_{\mathrm{x}}(\hat{N}(\mathscr{V}_{1})-\hat{N}(\mathscr{V}_{2}))|^{2}d\mathrm{x}$ (5.18)
$\leqslant\int_{\mathbb{R}^{d}}2|\mathscr{V}_{1}-\mathscr{V}_{2}|^{2}|\partial^{\gamma}_{\mathbf{x}}\big(\sum_{k=2}^{l_{0}}\alpha_{k}\sum_{i=2}^{k}\binom{k}{i}u_{0}^{k-i}\sum_{j=0}^{i-1}\mathscr{V}^{i-1-j}_{1}\mathscr{V}^{j}_{2}\big)|^{2}$
$\quad+2|\partial^{\gamma}_{\mathbf{x}}\mathscr{V}_{1}-\partial^{\gamma}_{\mathbf{x}}\mathscr{V}_{2}|^{2}|\sum_{k=2}^{l_{0}}\alpha_{k}\sum_{i=2}^{k}\binom{k}{i}u_{0}^{k-i}\sum_{j=0}^{i-1}\mathscr{V}^{i-1-j}_{1}\mathscr{V}^{j}_{2}|^{2}d\mathbf{x}$
[…]
Using (5.18) and summing over $\gamma$ with $0\leqslant|\gamma|\leqslant 1$ to (5.17), then applying Gronwall’s inequality and using the definition of $\|\cdot\|$ yields
[…]
where $C_{5}(T)$ has the same dependencies as $C_{4}(T)$ and is monotonically increasing in $T$ with $C_{5}(0)<\infty$. To ensure $\|\mathscr{W}\|=\|\mathscr{F}[\mathscr{V}_{1}]-\mathscr{F}[\mathscr{V}_{2}]\|\leqslant\frac{1}{2}\|\mathscr{V}_{1}-\mathscr{V}_{2}\|$, we require by (5.19),
$C_{5}(T)T\leqslant\frac{1}{4{\varepsilon}}.$ (5.20)
Set the lifespan time
$T_{\mathrm{lifespan}}:=\max\{T\mid C_{4}(T)T\leq\varepsilon^{-1/2}\ \text{and}\ C_{5}(T)T\leq(4\varepsilon)^{-1}\}.$ (5.21)
For such $T$ the map $\mathscr{F}$ is a contraction on $X_{\varepsilon}$, and Banach’s fixed point theorem gives a unique solution $\mathscr{U}\in X_{\varepsilon}$ to (5.1). Sobolev embedding provides the Hölder regularity.
Step 4: Uniform lower bounds for $T$ of problem (5.1). Define
$T_{\mathrm{lower}}:=\max\{T|C_{4}(T)T\leqslant 1\ \text{and}\ C_{5}(T)T\leqslant\frac{1}{4}\}.$ (5.22)
Since $\varepsilon<1$, we have $\frac{1}{\sqrt{\varepsilon}}>1$ and $\frac{1}{4\varepsilon}>\frac{1}{4}$. As $C_{4}(T)T$ and $C_{5}(T)T$ are continuous and increasing, it follows that $T_{\mathrm{lifespan}}\geqslant T_{\mathrm{lower}}$, with $T_{\mathrm{lower}}$ independent of $\varepsilon$ which implies that properties 3 holds.
The proof is complete. ∎
Remark 5.3.
The argument in the proof of Theorem 5.2 shares the same spirit as that of Theorem 3 in Section 12.2 of [7]. To ensure that the mapping $\mathscr{F}$ preserves the space $X_{\varepsilon}$ and satisfies the contraction property, the time $T$ is chosen to satisfy conditions (5.15) and (5.20). Notably, as $\varepsilon$ decreases, the existence time $T$ for problem (1.2) increases, thereby generalizing the conclusion of Theorem 3 in Section 12.2 of [7] in terms of the existence time $T$. In our setting, $T$ can be chosen arbitrarily large, whereas Theorem 3 in Section 12.2 of [7] generally guarantees only a short-time existence. This enhancement stems from three key structural features: (i) the problem (5.1) has trivial initial data; (ii) the polynomial nonlinearity $\hat{N}$ has compact support in $D$ and an explicit form; and (iii) the source term $F_{2}$ has an explicit expression, with its $H^{2}(D)$-norm being small in terms of $\varepsilon$. Combined with the careful construction of the closed set $X_{\varepsilon}$ and the properties of the function $u_{0}$, these features enable the existence of a solution to problem (5.1) over an arbitrarily large time interval $[0,T]$.
We now proceed to prove Theorem 1.2. As the argument closely parallels that of Theorem 1.1, we provide the complete details below for clarity and self-containedness.
Proof of Theorem 1.2..
In the three-dimensional case, according to Remark 3.8, we select the predetermined parameters $r_{0}$, $\varepsilon$, and $m$ as follows:
$\hfill r_{0}=\min\left\{\frac{1}{3}(\mathcal{M}+1)^{-1},\ \frac{1}{6}\min_{1\leq i<j\leq n}|\mathbf{x}_{i}-\mathbf{x}_{j}|\right\},$ (5.23)
$\hfill\varepsilon=\min\left\{\frac{1}{C_{1}+(C_{6})^{2}},\frac{1}{\left(C_{4}(T)T\right)^{2}},\frac{1}{4C_{5}(T)T},1\right\},$
$\hfill m=\max\left\{\left\lfloor\frac{256(2+C_{2}(r_{0})\varepsilon)^{2}r_{0}^{3}-3}{2}\right\rfloor+1,M(r_{0})\right\},$
where the constants $C_{1}$, $C_{2}(r_{0})$, and $M(r_{0},\omega)$ are specified in Theorem 3.1; Constants $C_{4}(T)$ and $C_{5}(T)$ are given by (5.7) and (5.19), respectively, in Theorem 5.2; Constant $C_{6}$ is given in Theorem 5.2; and the prescribed time $T$ is as given in Theorem 1.2. Then, we choose the initial inputs for problem (1.2) as specified in (3.22), with the corresponding $u_{0}$ determined by the parameters $r_{0}$, $\varepsilon$, and $m$ in (5.23). Let $\mathscr{U}$ denote the unique solution to auxiliary problem (5.1). A direct computation verifies that
$U:=u_{0}+\mathscr{U}$ (5.24)
constitutes the unique solution to problem (1.2). Since $u_{0}$ is smooth and defined on $0,\infty)$, the lifespan of $U$ is entirely determined by that of $\mathscr{U}$. As established by the lifespan of $\mathscr{U}$ in (5.21) and the explicit expression for $\varepsilon$ in (5.23), $U$ exists on the interval $[0,T]$, where $T$ denotes the prespecified time given in Theorem 1.2. Moreover, Theorems 5.2 and 3.1 together with Remark 2.3 yield $U\in L^{\infty}\big(0,T;H^{3}_{\text{loc}}(\mathbb{R}^{d})\big)$.
Next, applying the mean value theorem to (3.1) and (3.3) in Theorem 4.1 yields points $\D¯\mathbf{z}_{i}\in B_{3r_{0}}(\mathbf{x}_{i})\backslash\overline${D} ($i=1,2,\dots,n$) such that for all $t\in[0,T]$:
$|\nabla u_{0}(\mathbf{z}_{i},t)|\geqslant\frac{\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}-\left(C_{1}+C_{2}(r_{0})\right)\varepsilon}{3r_{0}}.$ (5.25)
Combining (5.23), (5.24), (5.25), and properties (3) in Theorem 5.2, we obtain
$|\nabla U(\mathbf{z}_{i},t)|$ $\geqslant\frac{\left(\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}-C_{2}(r_{0})\varepsilon\right)-C_{1}\varepsilon}{3r_{0}}-C_{6}\sqrt{\varepsilon}$ (5.26)
$>\frac{2-1}{3r_{0}}-1=\frac{1}{3r_{0}}-1>\mathcal{M},$
for a.e. $t\in[0,T]$. Here, we choose $m\geqslant\left\lfloor\frac{256(2+C_{2}(r_{0})\varepsilon)^{2}r_{0}^{3}-3}{2}\right\rfloor+1$ in (5.23) to guarantee that $\tfrac{1}{16}\sqrt{2m+3}\,r_{0}^{-3/2}-C_{2}(r_{0})\varepsilon>2$. On the other hand, using above estimate and properties (3) in Theorem 3.1, we calculate
[…]
for a.e. $t\in[0,T]$. By (5.26) and (5.27), properties (1) and (2) in Theorem 1.2 are verified.
Furthermore, under the above construction, the localized gradient amplification points are confined within the balls $B_{3r_{0}}(\mathbf{x}_{i})$ which implies that
[…]
as $\mathcal{M}\to\infty$, for almost every $t\in[0,T]$. Properties (3) in Theorem 1.2 is verified. The two-dimensional case follows analogously.
The proof is complete. ∎
Acknowledgment
The work of H. Diao is is supported by National Natural Science Foundation of China (No. 12371422), the Fundamental Research Funds for the Central Universities, JLU. The work of H. Liu is supported by the Hong Kong RGC General Research Funds (projects 11304224, 11311122 and 11303125), the NSFC/RGC Joint Research Fund (project N_CityU101/21), the France-Hong Kong ANR/RGC Joint Research Grant, A-CityU203/19.